OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Well, if lines are perpendicular, it means that there values are opposite recipricols of eachother. In general if your slope is m than the slope of a line perpendicular to it is -1/m. In this case the slope of the equation you are looking for is perpendicular to y=-3x+1 meaning the slope of the equation you are looking for is going to be 1/3 So, we have y=(1/3)x+b. To find b we can plug in the point that you already know. Since we know that the equation we are looking for passes through (-6,1) we can plug those numbers in. So: y=(1/3)x+b 1=(1/3)(-6)+b 1=-2+b 3=b so know we know the slope, and the y-intercept of your equation and that is enough to write the equation in slope intercept form. y=(1/3)x+3 do you understand the steps??